OpenStudy (anonymous):

Use the addition and subtraction formulas to find the exact value of the expression. Tan 165 degrees.

zepdrix (zepdrix):

Hmmm I think they probably want you to break it up like this. \[\large \tan(165^o)=\tan(135^o+30^o)\]

OpenStudy (anonymous):

what would you do next?

zepdrix (zepdrix):

\[\large \tan (a+b)=\frac{ \tan a + \tan b }{ 1- \tan a \tan b }\] Hmmm we can use the sum formula for tangent. 135 and 30 both correspond to special angles for the tangent, so it will break down pretty nicely.

zepdrix (zepdrix):

\[\large \tan(135+30)=\frac{ \tan135+\tan30 }{ 1-\tan135 * \tan30 }\] Understand the setup? :D

OpenStudy (anonymous):

well i tried it like this and it doesnt work out, or idk what to do next

OpenStudy (anonymous):

sorry its backwards! :o

zepdrix (zepdrix):

lolol backwards XD

OpenStudy (anonymous):

haah idk how to fix it :/

zepdrix (zepdrix):

Oh i see, you broke up the numbers differently. Looks like you were able to break it up so they're still special angles though, so that's good :) hmm

zepdrix (zepdrix):